OpenStudy (anonymous):

Find the midpoint of the line segment whose end points are given. (- 2, 2) and (3/2,4/3 )

OpenStudy (raden):

use the formula : xm = (x1 + x2)/2 ym = (y1 + y2)/2

OpenStudy (anonymous):

\[x(m)=-2+\frac{ 3 }{ 2 }\div 2\]

OpenStudy (raden):

give a bracket for (-2 + 3/2) / 2 = ...

OpenStudy (anonymous):

so do we cancel out the 2's

OpenStudy (raden):

maybe, will easier if u convert -2 to -4/2 so, (-2 + 3/2) / 2 = (-4/2 + 3/2) / 2 = (-1/2)/2 = -1/4

OpenStudy (raden):

do same idea, for the value of ym = (2 + 4/3)/2 simplify it

OpenStudy (anonymous):

so is the answer (5/3)

OpenStudy (raden):

yes correct... the mid point coordinat be (xm,ym) = (-1/4, 5/3)